An actor invests some money at 9% and $22000 more than four times the amount at 12%. The total annual interest earned from the investment is $44820. How much did he invest at each amount.

Answer by Boreal(15235)

x=9% amount
4x+22000=12%
They add to 44,820
.09x+.12(4x+22000)=44820
.09x+.48x+2640=44820
.57x=42180
x=$74000@9%=$6660
4x+22000=$318000@12% is $38160
They add to $44820.